OpenStudy (anonymous):

SOMEONE PLEASE HELP...Determine if triangle RST with coordinates R (3, 4), S (5, 5), and T (6, 1) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (mertsj):

Did you graph the triangle?

OpenStudy (anonymous):

how do i do that?

OpenStudy (solomonzelman):

I would maybe use a distance formula to find the sides, if you can't graph and say that it is because of what the graph looks like, that it is a right \(\large\color{black}{ \color{blue}{ \triangle }}\) or not a right \(\large\color{black}{ \color{blue}{ \triangle }}\).

OpenStudy (solomonzelman):

\(\large\color{black}{ D=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2}}\) (D is distance)

OpenStudy (anonymous):

im sorry i just havent done this since august.. and i have to do these 8 practice questions. but im not good at teaching myself. so it really sucks to be homeschooled.

OpenStudy (solomonzelman):

do you understand the distance formula?

OpenStudy (anonymous):

which one is that?..

OpenStudy (solomonzelman):

the one I posted above, \(\large\color{black}{ D=}\) etc.,

OpenStudy (anonymous):

well if i said yes i would be lying lol :(

OpenStudy (solomonzelman):

\(\large\color{black}{ D=\sqrt{(y_1-y_2)^2+(x_1-x_2)^2} }\) is the distance between the points, \(\large\color{black}{ (x_1,y_1) }\) and \(\large\color{black}{ (x_2,y_2) }\)

OpenStudy (anonymous):

okay

OpenStudy (solomonzelman):

For example, \(\large\color{black}{ D=\sqrt{(\color{blue}{4}-\color{magenta}{5})^2+(\color{red}{3}-\color{green}{5})^2} }\) is the distance between the points, \(\large\color{black}{ R~~ (\color{red}{3}, \color{blue}{4}) }\) and \(\large\color{black}{ S~~ (\color{green}{5}, \color{magenta}{5}) }\) .

OpenStudy (solomonzelman):

Do you see how I am plugging them in?

OpenStudy (anonymous):

those are the numbers from my question?

OpenStudy (solomonzelman):

yes they are.

OpenStudy (anonymous):

okay yeah i get that

OpenStudy (solomonzelman):

okay, tell me the exact value for the distance between R and S. Can you do that?

OpenStudy (solomonzelman):

you will just have to simplify the \(\large\color{black}{ D=\sqrt{(\color{blue}{4}-\color{magenta}{5})^2+(\color{red}{3}-\color{green}{5})^2} }\)

OpenStudy (solomonzelman):

simplify the right side....

OpenStudy (anonymous):

i did (4−5)2+(3−5)2 this and my calcultor got me -6?

OpenStudy (anonymous):

( 4 − 5 ) 2 + ( 3 − 5 ) 2 and got -6

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

(-1)^2 + (-2)^2 would be the first result after subtracting inside each of the parenthesis, correct?

OpenStudy (solomonzelman):

then (-1)^2=1 and (-2)^2=4, so the exact distance from S to R is?

OpenStudy (anonymous):

oh yeah i totally did that wrong..

OpenStudy (anonymous):

would the answer be a coordinate?

OpenStudy (solomonzelman):

no, it would be a value. A distance. (it would be a square rot of some number)

OpenStudy (anonymous):

wait so how would you even get a answer out of those?

OpenStudy (solomonzelman):

you should be able to just simplify inside the square root, reminding, it is For example, \(\large\color{black}{ D=\sqrt{(\color{blue}{4}-\color{magenta}{5})^2+(\color{red}{3}-\color{green}{5})^2} }\) and it is the distance between the 2 points, \(\large\color{black}{ R~~ (\color{red}{3}, \color{blue}{4}) }\) and \(\large\color{black}{ S~~ (\color{green}{5}, \color{magenta}{5}) }\) .

OpenStudy (mertsj):

Much easier if you just graph the points.

OpenStudy (mertsj):

Start at (0,0). To graph (3,4) go 3 units to the right and 4 units up. Make a dot there.

OpenStudy (anonymous):

is there site i could use to graph stuff?

OpenStudy (mertsj):

To graph (5,5) go 5 units to the right and 5 units up.

OpenStudy (mertsj):

To graph (6,1) go 6 units to the right and 1 unit up.

OpenStudy (anonymous):

ik how to graph stuff. but i dont have graph paper. so is there something online i could use?

OpenStudy (mertsj):